OpenStudy (anonymous):

the terminal side of x=23pi/3 lies in (a) quad 1 (b) quad 2 (c) quad 3 (d) quad 4

OpenStudy (mertsj):

\[\frac{23\pi}{3}=7 \frac{2}{3}\pi \]

OpenStudy (mertsj):



OpenStudy (mertsj):